OpenStudy (anonymous):

A sandbag was thrown downward from a building. The function f(t) = -16t2 - 32t + 128 shows the height f(t), in feet, of the sandbag after t seconds. Part B: Complete the square of the expression for f(x) to determine the vertex of the graph of f(x). Would this be a maximum or minimum on the graph? Part C: Use your answer in part B to determine the axis of symmetry for f(x)?

OpenStudy (anonymous):

@mathmale @mathstudent55 @satellite73 @triciaal @ganeshie8 @sleepyjess

OpenStudy (anonymous):

you do not (repeat do NOT) need to complete the square to do this

OpenStudy (anonymous):

How do you complete a square?

OpenStudy (anonymous):

the first coordinate of the vertex of \(y=ax^2+bx+c\) is always \[x=-\frac{b}{2a}\]

OpenStudy (anonymous):

@satellite73 Perhaps, but the question is asking me to complete the square nonetheless.

OpenStudy (anonymous):

the second coordinate is what you get when you replace \(x\) by the first coordinate

OpenStudy (anonymous):

in your case \(a=-16,b=-32\)

OpenStudy (anonymous):

And I already know how to calculate the vertex, thanks! But how do I complete a square @satellite73 ? (This is unrelated, but how did you get a 100 smartscore, and are you the only person to ever get a 100 smartscore?)

OpenStudy (anonymous):

ok first off lets note that the question makes no sense at all physically, since \(t\) represents time, and the first coordinate of the vertex is going to be negative so unless you can go backwards in time, (which you cannot) the question is meaningless

OpenStudy (anonymous):

Yeah, I pointed that out in Part A

OpenStudy (anonymous):

But anyways, how do you complete a square?

OpenStudy (anonymous):

Theoretically you could go "forward in time" by using a black hole's gravitational pull, but that's unrelated.

OpenStudy (anonymous):

think about this: you throw a rock DOWN from an initial height, and you want to know the maximum height pretty clearly the max is where you started from

OpenStudy (anonymous):

Yeah I get it, the question doesn't make much sense. But how do you complete a square anyways?

OpenStudy (anonymous):

if you still want to complete the square, we can do it

OpenStudy (anonymous):

Yes please

OpenStudy (anonymous):

first off, factor out the \(-16\) from the first two terms what do you get ?

OpenStudy (anonymous):

You can factor out -16 from all 3 terms, shouldn't I do that?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

k f(t) = -16(t2 - 2t) + 128

OpenStudy (anonymous):

ok so we are at \[f(t)=-16(t^2-2t)+128\]now to complete the square, what is half of \(-2\)?

OpenStudy (anonymous):

-1?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and \((-1)^2=?\)

OpenStudy (anonymous):

-1?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

oh wait, sorry. 1

OpenStudy (anonymous):

eys

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

ok that was wrong, let me back up

OpenStudy (anonymous):

we got \((-1)^2=1\) right?

OpenStudy (anonymous):

wait, it was 2t, where did the t go?

OpenStudy (anonymous):

lets ignore the \(-16\) for the moment, and concentrate only on \(t^2-2t\)

OpenStudy (anonymous):

right

OpenStudy (anonymous):

but how did you make the t go away?

OpenStudy (anonymous):

half of \(-2\) is \(-1\) so we are going to complete the square by writing \((t-1)^2\)

OpenStudy (anonymous):

now the \(t\) did not go away, because \[(t-1)^2=t^2-2t+1\] right?

OpenStudy (anonymous):

but something has changed we started with \[t^2-2t\] and ended with \[t^2-2t+1\] so we completed the square by adding \(1\)

OpenStudy (anonymous):

I dont really get it...

OpenStudy (anonymous):

when you complete the square, you are making a change to write it as a perfect square

OpenStudy (anonymous):

(t−1)2=t2−2t+1

OpenStudy (anonymous):

we have to compensate for that change

OpenStudy (anonymous):

oh, because you divided 2 in half, so you added 1 somewhere else?

OpenStudy (anonymous):

right, add or subtract as the case may be

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

now is where it gets a little complicated so lets go slow

OpenStudy (anonymous):

k

OpenStudy (anonymous):

we have \[-16(t^2-2t)+128\] and we write the first term as a perfect square as \[-16(t-1)^2+128\pm?\]

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

now the first term is a perfect square, but we have to figure out what change we made so we can compensate for it

OpenStudy (anonymous):

and compensate for it next to 128?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

i think we only took away 1. That's the only change i think, correct?

OpenStudy (anonymous):

yes, but we actually have to worry about the \(-16\) in front

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

we have also made a mistake i see, so lets correct it first

OpenStudy (anonymous):

wheres the mistake?

OpenStudy (anonymous):

\[-16t^2-32t+128=-16(t^2+2t)+128\]

OpenStudy (anonymous):

i don't see any mistake

OpenStudy (anonymous):

the process is still the same, half of 2 is 1, one squared is 1

OpenStudy (anonymous):

we had \(-2t\) when it should have been \(+2t\)

OpenStudy (anonymous):

oh yeah

OpenStudy (anonymous):

so now we are at \[-16(t+1)^2+128\pm?\]

OpenStudy (anonymous):

why is it 128 +-?

OpenStudy (anonymous):

now we didn't really add 1, because of that \(-16\) out fron

OpenStudy (anonymous):

that is what we are going to correct

OpenStudy (anonymous):

we have to figure out what we added or subtracted so we can compensate for it

OpenStudy (anonymous):

wait, how did we not really add 1. Okay, this is the most complicated math problem i've ever had

OpenStudy (anonymous):

gaah

OpenStudy (anonymous):

How does that -16 affect us not actually adding a 1

OpenStudy (anonymous):

it looks like we added one (don't fret, i am going to show you a real snap way to do this )

OpenStudy (anonymous):

I'm confused :(

OpenStudy (anonymous):

\[-16(t+1)^2=-16(t^2+2t+1)=-16t^2-32t-16\]

OpenStudy (anonymous):

is that part clear ?

OpenStudy (anonymous):

It might help if you reversed the order, but yeah, I think i get that part

OpenStudy (anonymous):

ok so we actually did not add one, the -16 out front means we SUBTRACTED 16

OpenStudy (anonymous):

...

OpenStudy (anonymous):

you see it right? the \(-16\) out at the end there

OpenStudy (anonymous):

my expression isn't too far off from my profile pic right now

OpenStudy (anonymous):

cause I really have no idea what you just said, or how you got to that answer @satellite73

OpenStudy (anonymous):

(This is unrelated, but how did you get a 100 smartscore, and are you the only person to ever get a 100 smartscore?)

OpenStudy (anonymous):

\[-16(t+1)^2=-16(t^2+2t+1)=-16t^2-32t\color{red}{-16}\]

OpenStudy (anonymous):

you do see it right? it is red

OpenStudy (anonymous):

yeah, i see the -16, but how does that mean we never added a 1?

OpenStudy (anonymous):

we added 1 to complete the square, but because of the -16 out front we really subtracted 16

OpenStudy (anonymous):

so we have to add it back

OpenStudy (anonymous):

WAIT HOLD ON

OpenStudy (anonymous):

\[-16t^2-32t+128=-16(t+1)^2+128+16\] that equal sign is for real

OpenStudy (anonymous):

how does it mean you actually subtracted 16? I really don't get this

OpenStudy (anonymous):

Ya know, I might be able to complete this question by saying you don't need to find the square to find the vertex...

OpenStudy (anonymous):

we changed \(t^2+2t\) in to \((t+1)^2\) by adding 1 right, because \((t+1)^2=t^2+2t+1\) is that clear or no?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

And then you would compensate by adding 1 somewhere else

OpenStudy (anonymous):

yes, if that was all we had, that is what we would do (subtract 1 to compensate, not add)

OpenStudy (anonymous):

we add one, we have to subtract one